OpenStudy (anonymous):

v^-0.84=0.0718325792, solve for v.

jimthompson5910 (jim_thompson5910):

hint: \[\Large x^k = y\] \[\Large (x^k)^{1/k} = y^{1/k}\] \[\Large x^{k*1/k} = y^{1/k}\] \[\Large x^{k/k} = y^{1/k}\] \[\Large x^{1} = y^{1/k}\] \[\Large x = y^{1/k}\]